If the ΔH°rx and ΔS°rx for a gas phase chemical reaction are 419.53 kJ and -178.85 J/K respectively, determine ΔG°rx (in kJ) at 103.11 °C. Report your answer to two decimal places (i.e. 123.45 kJ). Assume that ΔH°rx and ΔS°rx do not vary with temperature.
Answer #1

Apply:

G = H - TS

G = (419.53*1000) - (103.11+273)*(-178.85) = 486797.2735 J/mol

G = 48.68 kJ/mol